A new business requires $100,000 investment. The project can either be entirely equity financed, or 70% of the investment can be funded by a bank loan at 10%. Assume no tax. EBIT is expected to be $25,000, but it could be as low as $3,000.

b) At what level of EBIT the shareholders receive the same return on equity regardless the percentage of debt financing?
Answer #1

Solution:

Given data,

A new business requires $100,000 investment. The project can either be entirely equity financed, or 70% of the investment can be funded by a bank loan at 10%. EBIT is expected to be $25,000, but it could be as low as $3,000.

Now solution:

Return on Equity = Earnings for Equity / Equity Funds

Take EBIT as A.

Then

A / 100,000 = (A - 7,000) / 30,000

0.3A = A - 7,000

A = $10,000.

When EBIT = 30,000,The shareholders will receive the same return on equity regardless of the percentage of debt financing.